OpenStudy (anonymous):

find the x- and y- intercepts -4x+2y=16

OpenStudy (aroub):

If you want to find the x-intercept you substitute y by 0 and when you want to find the y-intercept you substitute the x by 0

OpenStudy (anonymous):

x-intercept ( 0, 8) y -intercept ( -4, 0)

OpenStudy (aroub):

so you solve it like this : -4x+2y=16 the x-intercept: -4x+2(0)=16 -4x=16 x=-4 (-4,0) the y-intercept: -4(0)+2y=16 2y=16 y=8 (0,8)

OpenStudy (aroub):

Hmm, it's the other way around chlorophyll ?